OpenStudy (anonymous):

A medical screening program administers three independent fitness tests. Of the persons taking the tests, 85% pass test 1, 65% pass test 2, and 50% pass test 3. A participant is chosen at random a) What is the probability that she will pass all 3 tests? b) What is the probability that she will pass at least 2 of the 3 tests?

OpenStudy (anonymous):

prob of passing test 1 = 0.85 prob of passing test 2 = 0.65 prob of passing test 3 = 0.65

OpenStudy (anonymous):

here the events are all independent so just multiply the individual probabs for part (a)

OpenStudy (anonymous):

multiply all 3 together?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

okay i got .27625 for that!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how do i find the second?

OpenStudy (anonymous):

can you say the probab of not passing test 1

OpenStudy (anonymous):

prob of passing test 1 = 0.85 hence prob of not passing test 1 is 1-0.85 =0.15 and similar for the rest tests

OpenStudy (anonymous):

so do i find all of those and use them together?

OpenStudy (anonymous):

u need to find them but also check what the question asks

OpenStudy (anonymous):

hint : prob for passing test 2 and test3 and not tets1 = 0.65*0.5*0.15 =0.04875

OpenStudy (anonymous):

1-0.04875?

OpenStudy (anonymous):

nope find the probab for apssing test1 and 3 and not test2 find the probab for apssing test1 and 2 and not test3

OpenStudy (anonymous):

ahh okay hold on

OpenStudy (anonymous):

not passing test 2 would be .35*.85*.5= .14875 not passing test 3 would be .5*.85*.65= .27625

OpenStudy (anonymous):

probability that she will pass at least 2 of the 3 tests = (prob for passing test 2 and test3 and not tets1) + (probab for apssing test1 and 3 and not test2)+ (probab for apssing test1 and 2 and not test3) + probability that she will pass all 3 tests

OpenStudy (anonymous):

.75?

OpenStudy (anonymous):

?